Prev Next

SW&IT Meeting Minutes 2019-12-02

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191104
LR20191104 - Bec, Matthieu 80.0
LR20191104 - Holmberg, Carl 80.0
LR20191104 - Colson, Andrew 56.0
LR20191104 - Edmeades, David 48.0
LR20191104 - Sola, Igor 63.0
LR20191104 - Kubanek, Petr 100.7
LR20191104 - Hooper, Stephen 80.0
LR20191104 - Zhang, Yang 70.0
LR20191104 - Zhang, Xianyu 72.0

20191118
LR20191118 - Bec, Matthieu 64.0
LR20191118 - Holmberg, Carl 64.0
LR20191118 - Colson, Andrew 48.0
LR20191118 - Edmeades, David 64.0
LR20191118 - Sola, Igor 40.0
LR20191118 - Kubanek, Petr 24.0
LR20191118 - Hooper, Stephen 64.0
LR20191118 - Zhang, Yang 61.0
LR20191118 - Zhang, Xianyu 40.0

News from management

Previous Action Items

  • web technologies
    • currently
      • IMS/DMS (polymer, wsgi)
      • statserv, info (php ?)
      • mini-q (react, )
    • convergence

Open topics

  • Follow up: support schedule, liason, response, day/night time
  • TMS reschedule tests

Team reports

MB

  • OC, WPC, Management
  • Meeting on shutdown/restart and documenting procedure
  • DMS zabbix integrated, but some WIP
  • AO data meeting
next
  • Obs WS presentaton to sciops
  • Power restart document, inc. loss of cooling
  • DMS hashed urls, issue at startup
  • sql zabbix closer inspection
  • AO data follow up
  • chksum header check with sciarc
  • LBTI alarm meeting

DE

  • Purchases meeting
    • Are we cleared for the whole list?
  • VM spin-ups for Yang and Carl
  • TMS Win7 machine space cleanup
  • All drives in stock for mountain Synology volume
  • New safety training
  • Projector install
    • JH immediately remarked on the presence of reds
  • IT email list setup for CB

CH

Last week
  • ALH: Draft state diagram for LBTI integration.
  • ARGOS: Reviewed design docs, discussed build process with AC, requested dev vm.
  • Mgmt: All hands meeting.
  • Supportscripts: Added patch to restart DS9 after crash.
  • TMS: Updated config files. TMS host/laser test detects state, but doesn't update properties properly.

This week
  • ALH: Prototype an LBTI agent.
  • ARGOS: Run build process for Arbitrator, controllers.
  • Supportscripts: Retry DS9 restart patch test. Last weekend's login dropped out.
  • TMS: Test status script/ALH integration.
  • Safety training.

YZ

  • Working on MODS/LBC code and program setup for 64-bit machines.
  • Continued discussion for IT#7927 (MODS image corrupted during transfer).
  • Update ModsView code on Github with Olga.
  • AO-Cravat database rebuild and interface redesign.

AC

  • ARGOS
    • RPM  - "missing" script is too old or missing
    • check state of software deployed on various argos machines and sync as necessary
    • arbitrator simulator and software installation overview with CH
    • test lite observer status GUI

IS

Last week:
  • PIT
    • worked on ways to improve and automatize proposals after period closes.
    • worked on implementing a resubmit option to the PIT so there are no or to limit duplicates
This week
  • PIT
    • ongoing work on resubmission logic
  • miniQ
    • clean up and make ingesting or setting up new run, easier (only from backend)
  • LUCI
    • Need to do some code repository cleaning

XZ

Last week:
  • AO data document
  • AO data meeting
  • Discussion with SOUL people
  • Document MOXA safe and power on status
  • define steps for IDL license serverce replacement
This week:
  • Setup the mountain anallyze VM
  • 4D test support
  • AO plan eval and Est meeting
  • AO telemetry document
  • SOUL lab setup support?

SH

  • PEPSI sorted
  • Ordering equipment

next:

  • ovirt
  • docker repository
  • finish slides for obs workstation upgrade
  • re-home vm1.tucson
Topic revision: r13 - 02 Dec 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback